# Structure-Based Molecule Optimization via Gradient-Guided Bayesian Update

### v2

@@ Appendix G Evaluation of Molecular Conformation
To measure the quality of generated ligand poses, we further employ PoseCheck Harris et al. (2023) to calculate the Strain Energy (Energy) of molecular conformations and Steric Clashes (Clash) w.r.t. the protein atoms in Figure 10 and 11, respectively.
@@
We report the ratio of redocking RMSD below 2Å between generated poses and Vina docked poses to reveal the agreement of binding mode. Due to issues of poses generate by Autodock, not all pose pairs are available for calculating symmetry-corrected RMSD, where we report the non-corrected RMSD instead to make sure that all samples are faithfully evaluated. As shown in Figure 12, the optimization methods all display a tendency towards generating a few outliers, which might be attributed to the somewhat out-of-distribution (OOD) nature of optimization that seeks to shift the original distribution. Among all, DecompOpt generates the most severe outliers with RMSD as high as 160.7 Å, and its unsatifactory performance is also suggested by the lowest ratio of RMSD $<$ 2Å (24.3%), while for gradient-based TAGMol and our method, it only has a negligible impact and the ratio is generally more favorable.
@@
Energy Med. ($\downarrow$) Clash Avg. ($\downarrow$) RMSD $<$ 2Å ($\uparrow$) Connected ($\uparrow$) [^1] Connected ratio of DecompOpt and RGA is calculated based on the optimization results of all rounds provided by the authors. For each of the $N$ rounds, a total of $k$ molecules ought to be generated, thus we divide the total number of optimized molecules for n pockets by $N\times k\times n$.
@@ Appendix H Inference Time
Table 12: Inference time cost of optimization baselines.
@@ Appendix I More Related Works
Molecule Optimization
As an alternative to target-aware generative modeling of 3D molecules, the optimization methods are goal-directed, obtain desired ligands usually by searching in the drug-like chemical space guided by property signals (Bilodeau et al., 2022; Du et al., 2024). General optimization algorithms were originally designed for ligand-based drug design (LBDD) and optimize common molecule-specific properties such as LogP and QED (Olivecrona et al., 2017; Jin et al., 2018; Nigam et al., 2020; Spiegel & Durrant, 2020; Xie et al., 2021; Bengio et al., 2021), but could be extended to structure-based drug design (SBDD) given docking oracles. However, since most early attempts did not take protein structures into consideration thus were essentially not target-aware, it means that they need to be separately trained on the fly for each protein target when applied to pocket-specific scenarios. RGA (Fu et al., 2022) explicitly models the protein pocket in the design process, overcoming the transferability problem of previous methods.
Appendix J Limitation
Limitation of this work lies in that as the first proof-of-concept, MolJO adopts no more than three objectives (Affinity, QED, SA) in the optimization process, while there are more objectives that are biologically meaningful such as ADMET. Future directions include extending the general framework to accommodate a wider range of objectives, and even beyond the scope of structure-based molecule optimization (SBMO), for MolJO is a general gradient-based optimation method for continuous and discrete variables, and can be tailored to a great many read-world applications.

### v3

@@ Appendix G Evaluation of Molecular Conformation
To measure the quality of generated ligand poses, we further employ PoseCheck (Harris et al., 2023) to calculate the Strain Energy (Energy) of molecular conformations and Steric Clashes (Clash) w.r.t. the protein atoms in Fig. 9 and 10, respectively.
@@
We report the ratio of redocking RMSD below 2Å between generated poses and Vina docked poses to reveal the agreement of binding mode. Due to issues of poses generate by Autodock, not all pose pairs are available for calculating symmetry-corrected RMSD, where we report the non-corrected RMSD instead to make sure that all samples are faithfully evaluated. As shown in Fig. 11, the optimization methods all display a tendency towards generating a few outliers, which might be attributed to the somewhat out-of-distribution (OOD) nature of optimization that seeks to shift the original distribution. Among all, DecompOpt generates the most severe outliers with RMSD as high as 160.7 Å, and its unsatifactory performance is also suggested by the lowest ratio of RMSD $<$ 2Å (24.3%), while for gradient-based TAGMol and our method, it only has a negligible impact and the ratio is generally more favorable.
@@
Energy Med. ($\downarrow$) Clash Avg. ($\downarrow$) RMSD $<$ 2Å ($\uparrow$) Connected ($\uparrow$) [^1] Connected ratio of DecompOpt and RGA is calculated based on the optimization results of all rounds provided by the authors. For each of the $N$ rounds, a total of $k$ molecules ought to be generated, thus we divide the total number by $N\times k\times n$.
@@ Appendix H Inference Time
Table 16: Inference time cost of optimization baselines, error bars indicating the standard deviation across 10 randomly selected proteins.
@@ Appendix I More Related Works
General Molecule Optimization.
As an alternative to target-aware generative modeling of 3D molecules, the optimization methods are goal-directed, obtain desired ligands usually by searching in the drug-like chemical space guided by property signals (Bilodeau et al., 2022; Sun et al., 2023; Du et al., 2024). General optimization algorithms were originally designed for ligand-based drug design (LBDD) and optimize common molecule-specific properties such as LogP and QED (Olivecrona et al., 2017; Jin et al., 2018; Nigam et al., 2020; Spiegel & Durrant, 2020; Xie et al., 2021; Bengio et al., 2021), but could be extended to structure-based drug design (SBDD) given docking oracles. However, since most early attempts did not take protein structures into consideration thus were essentially not target-aware, it means that they need to be separately trained on the fly for each protein target when applied to pocket-specific scenarios. RGA (Fu et al., 2022) explicitly models the protein pocket in the design process, overcoming the transferability problem of previous methods. DiffAC (Zhou et al., 2024b) utilizes policy gradients for SDEs to fine-tune pretrained diffusion models given affinity signal, demonstrating the potential of RL method but limited to Vina Score optimization only.
Constraint Molecule Optimization.
Real-world lead optimization typically requires retaining specific substructures to preserve critical molecular interactions or properties. Recently, CBGBench (Lin et al., 2025) introduces a generative graph completion framework for systematic evaluation, including tasks such as linker design, fragment growing, side chain decoration, and scaffold hopping, showing the potential for rigorous benchmarking of structure-based molecular optimization under predefined substructural constraints.